# Ab initio correlated calculations without finite basis-set error: Numerically precise all-electron RPA correlation energies for diatomic molecules

### v1

@@ Abstract
MAE / / 0.07
In Table S4 of the SM, we further present the FC-RPA binding energies for a larger set of diatomic molecules, obtained using the Sternheimer approach and conventional schemes with the largest available GTO cc-pV6Z/cc-pV5Z basis set, one still has an MAE of 1.60kcal/mol. Fortunately, by extrapolating to the CBS limit, the BSIE can be significantly reduced to below 0.4 kcal/mol, for both GTO and NAO basis sets. However, for heavy elements, such correlation-consistent basis sets are unavailable, making conventional approaches unusable to obtain high-quality RPA results. In contrast, our basis-error-free RPA results provide invaluable references for developing high-quality AO basis sets for heavy elements. This possibility will be explored in future work.
We note that the numerical technique developed in the present work for correlated calculations without basis errors does not only apply to RPA, but is also directly extendable to any method based on density response functions $\chi^{0}$. One prominent example is the SOS-MP2 approach [29, 30], which can be calculated as

### v2

@@ Abstract
MAE / / 0.05
In Table S4 of the SM, we further present the FC-RPA binding energies for a larger set of diatomic molecules, obtained using the Sternheimer approach and conventional schemes with the largest available GTO cc-pV6Z/cc-pV5Z basis set, one still has an MAE of 1.60 kcal/mol. Fortunately, by extrapolating to the CBS limit, the BSIE can be significantly reduced to below 0.4 kcal/mol, for both GTO and NAO basis sets.However, for heavy elements, such correlation-consistent NAO basis sets are unavailable, while the performance of currently available GTO basis sets developed for heavy elements for RPA calculations remains largely unexplored. In contrast, our basis-error-free RPA results provide invaluable references for developing high-quality AO basis sets for heavy elements. This possibility will be explored in future work.
We note that the numerical technique developed in the present work for correlated calculations without basis errors not only applies to RPA, but is also directly extendable to any method based on density response functions $\chi^{0}$. One prominent example is the SOS-MP2 approach [37, 38], which can be calculated as